OpenStudy (anonymous):

Solve and check. 4^2x= 32^1/2 Can someone walk me through the steps? Thank you!! :)

OpenStudy (anonymous):

32 to the power of 1/2

OpenStudy (anonymous):

@mthompson440 are you trying to simplify both sides of the equation?

OpenStudy (anonymous):

I don't know..?

OpenStudy (anonymous):

32^1/2= 16

OpenStudy (anonymous):

Divide both sides by 16.

OpenStudy (jhannybean):

\[\large4^{2x}=32^{1/2}\]\[\large (2^2)^{2x} = (2^5)^{1/2}\]\[\large 2^{4x} = 2^{5/2}\]\[\large 4x = \frac{5}{2}\]\[\large x=~?\]

OpenStudy (anonymous):

Ok.. But I don't understand how to solve it.

OpenStudy (anonymous):

@Jhannybean explained it

OpenStudy (jhannybean):

Your goal is to make the same base so you can evaluate their powers.

OpenStudy (anonymous):

X= 5/8?

OpenStudy (anonymous):

incorrect

OpenStudy (jhannybean):

32 = 2 * 2 * 2 * 2 * 2 4 = 2 * 2 Theres an exponent rule that states having the same base allows for you to evaluate powers. \[\large a^x = a^y \implies x=y\]

OpenStudy (jhannybean):

Yes @calecea . That is correct.

OpenStudy (anonymous):

@Jhannybean i thought it was 1

OpenStudy (anonymous):

@mthompson440 I kind of get it. I mean it's half of 32, but it was just confusing it's it's the exponent, not like multiplication.

OpenStudy (jhannybean):

\[4x=\frac{5}{2} \implies x = \frac{5}{2\cdot 4} = \frac{5}{8} \ne 1\]

OpenStudy (anonymous):

@jhannybean thank you!

OpenStudy (anonymous):

@Jhannybean Oh lol

OpenStudy (jhannybean):

@mthompson440 32\(^{1/2}\) \(\ne\) 16

OpenStudy (anonymous):

How do you get the 1/2 up there like that?

OpenStudy (jhannybean):

\[32^{1/2} =\sqrt{32} = \sqrt{16 \cdot 2} = \sqrt{4^2 \cdot 2} = 4\sqrt{2} \ne 16\]

OpenStudy (anonymous):

what do you mean, "up there like that?"

OpenStudy (anonymous):

like by the 32

OpenStudy (anonymous):

\[32^1/2\]

OpenStudy (jhannybean):

`^{1/2}`

OpenStudy (anonymous):

nvm figured it out

OpenStudy (jhannybean):

Have you learned exponential functions, @calecea ?

OpenStudy (jhannybean):

example problem: \(\large 7^{5x+3} = 512\)

OpenStudy (jhannybean):

Now try solving this one.

OpenStudy (anonymous):

I have no idea. Probably.

OpenStudy (jhannybean):

Using our knowledge of log rules, we know that: \(\log (a)^b \implies b\log(a)\)

OpenStudy (jhannybean):

Okay, let's try that one instead.

OpenStudy (jhannybean):

\[\large \log(x) +\log(x+48)=2\]

OpenStudy (jhannybean):

Let me know how far you've gotten with it.